GFTOPK(1) GFTOPK(1) NAME gftopk−convertgenericfontfilestopackedfontfiles SYNOPSIS gftopk[−v]gffilename[pkfilename] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thegftopkprogramconvertsagenericfontfileoutputby,forexample,mf(1),toapackedfont fileforusebyDVI-readingprograms.Packedfontfiles(pkfiles)aremuchsmallerthanthecorre-spondinggffiles,sotheyaregenerallythefontformatofchoice. Thegffilenameonthecommandlinemustbecomplete.Becausetheresolutionispartofthe extension,itwouldnotmakesensetoappendadefaultextensionasisdonewithTEXorDVIreadingsoftware.Thepkfilenamedefaultstothesame(stripped)nameasgffilename,anditis placedinthecurrentworkingdirectorywiththepksuffixreplacinggf.Forexample,theinputfile cmr10.300gfwouldbecomecmr10.300pk. OPTIONS Unlessthe−vswitchisgiven,gftopkoperatessilently.With−v,theversionnumberoftheprogramandstatisticsaboutthepackingareoutputtostdout. ENVIRONMENT gftopklooksforgffilenamefirstinthecurrentdirectory.Ifitisnotpresentthere,itusesthe environmentvariableGFFONTS.Ifthatisnotset,itusesthevariableTEXFONTS.Ifthatis notset,itusesthesystemdefault. Seetex(1)forthedetailsofthesearching. SEEALSO gftype(1),pktogf(1),pktype(1). DonaldE.Knuthetal.,METAFONTware. AUTHORS TomasRokickiwrotetheprogram.ItwaspublishedaspartoftheMETAFONTwaretechnical report,availablefromtheTEXUsersGroup.PaulRichardsoriginallyportedittoUnix. Web2C7.32x 14December1993 1